jabez177 (jabez177):

May someone please check my answer? :) What is the slope-intercept form of the function that contains the point (1, –2) and has a slope of –3? y = x +

jabez177 (jabez177):

I know first off the equation will be y=-3x+b because -3 is the slope and I think the y-intercept is -5 so my answer is \[y=-3x +-5\]

jabez177 (jabez177):

@imqwerty @Hayhayz

Hayhayz (hayhayz):

Why did you tag me xDDD

jabez177 (jabez177):

I thought you'd know these stuff...

imqwerty (imqwerty):

no thats not correct how did you take c as -5?

jabez177 (jabez177):

I used the Formula: y-y1=m(x-x1)

imqwerty (imqwerty):

yes thats correct :) just check your calculations once again

jabez177 (jabez177):

\[y-(-2)=-3(x-1)\] Did I put them in correctly?

imqwerty (imqwerty):

yes that is all correct continue

jabez177 (jabez177):

Now I add 2 to both sides, right?

imqwerty (imqwerty):

remember that negative times negative becomes positive so on the left when you multiply that -2 with outside negative sign you get a positive 2

jabez177 (jabez177):

I think that's where I messed up... May you write the equation now with the Positive 2? I lost track when you just said right now what you said.

imqwerty (imqwerty):

yes :) \(y \color{red}- (\color{red}- 2)=-3(x-1)\) \(y+2=-3(x)-3(-1)\)

jabez177 (jabez177):

Is my answer \[y=-3x+-1\]

jabez177 (jabez177):

Because I made the Negative twos positive.

imqwerty (imqwerty):

its not correct theres a little mistake u will be getting this-> \(y+2=-3x \color{red} +3\)

jabez177 (jabez177):

So my answer is \[y=-3x+5\]

imqwerty (imqwerty):

no after this equation-> \(y+2=-3x \color{red} +3\) you have to subtract 2 from both sides

jabez177 (jabez177):

So it is \[y=-3x-1~or~plus~1?\]

imqwerty (imqwerty):

\(y+2=-3x \color{red} +3\) \(y+2-2=-3x \color{red} +3-2\) \(y=-3x \color{red} +1\)

jabez177 (jabez177):

Thought it was plus 1

imqwerty (imqwerty):

yes it is +1

jabez177 (jabez177):

Okay, thank you, @Imqwerty! You're the best! :)

imqwerty (imqwerty):

lol ;)) np<3

jabez177 (jabez177):

Can you check this one? When 40 is added to the number of miles Karen ran last week, the result is the same as adding 10 to 4 times the number of miles she ran last week. How many miles did Karen run last week? mi

jabez177 (jabez177):

I think it's 10 miles

imqwerty (imqwerty):

yes correct :)

jabez177 (jabez177):

Tickets to a movie cost $9 for each adult and $4.50 for each child. Molly purchased 11 tickets for $72. How many adult and how many child tickets did she purchase? Molly purchased _ adult tickets and _ child tickets.

jabez177 (jabez177):

She purchased 5 adult tickets and 6 child tickets, right?

imqwerty (imqwerty):

correct :)

jabez177 (jabez177):

Thanks! :)

imqwerty (imqwerty):

anytime dude :)

jabez177 (jabez177):

I have 4 more questions for my homework, think you can check em after I'm done?

imqwerty (imqwerty):

ofc :)